Comparing PPP based price versus actual price of the latest iPhone in Melbourne, Australia and New York, given below:

Current Price of iPhone in US = $850

Current Price of iPhone in Aus = A$1150

Spot Rate (A$/$) = 1.3529

Annual Inflation rate in US = 1.50%

Annual Inflation rate in Aus = 3.35%

If prices were to remain steady and PPP were to hold, what should be the price in Melbourne, one year from now?
Answer #1

Spot rate (A$/$) = 1.3529

After one year from this spot rate = Future rate = Spot rate * ( 1 + Inflation rate in Aus) / ( 1 + Inflation rate in US)

= 1.3529 * (1.0335) / (1.0150)

= 1.3776

If prices of iPhone remain steady, after one year also we have the price of iPhone = $ 850 ( in US - place of origin)

But in Australia due to change in exchange rate, price of iPhone shall be = 850 * 1.3776 = 1170.96 ...... or A$ 1,171

Explanation

iPhone place of manufacture is US. Hence price remains steady implies that price remain same in US. But Australia had to import iPhone. As such price payable in terms of Australian dollars increases or decreases based on existing exchange rate at that time.

We see that Australian dollar to US dollar changed from A$ 1.3529 / $ to A$ 1.3776 after one year. Hence price of iPhone in Australia is likely to be 850 * 1.3776 = A$ 1,171.